OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality 2x < -10 or x + 3>(or equal to) 11 A number line with a closed circle on -5, shading to the left, and an open circle on 8, shading to the right. A number line with an open circle on -5, a closed circle on 8, and shading in between. A number line with an open circle on -5, shading to the left, and a closed circle on 8, shading to the right. A number line with a closed circle on -5, an open circle on 8, and shading in between.

OpenStudy (nory):

First, "solve" each inequality by isolating x on one side. Do this for both inequalities. What do you get?

OpenStudy (anonymous):

I'm really confuse in this type of math! :/

OpenStudy (anonymous):

confused*

OpenStudy (nory):

Remember how to solve a simple equation? Divide both sides of the first inequality by 2 to get x<-5. Then subtract three from both sides of the second to get x > 8.

OpenStudy (nory):

Can you plot these on the number line?

OpenStudy (anonymous):

ahh! I don't get it :(

OpenStudy (nory):

Which step didn't you understand?

OpenStudy (anonymous):

none, can you go threw it step by step?

OpenStudy (nory):

Okay. First you need your inequalities nice and simple. Your first one is 2x<-10. How do you simplify? Divide both sides by 2, just like an equation. Now what do you get?

OpenStudy (e.mccormick):

A note for doing this in general... FYI: I will use example numbers, and not the ones in your problem. "Solving" an inequality is the same as doing equalities, but with one small caveat. If you multiply or divide through with a negative number, it changes the direction of the inequality. For example, if I take \(a\le b\) and multiply through by \(-1\) it becomes \(-a\ge -b\) I do not see that happening in this problem, but it is something to be aware of when doing these. Other than the times when that happens, it is exactly like what Nory is saying. This is an equation you solve. Just like when they give you \(2a+1=3\) and say, "Solve for a." You do the whole \(2a+1-1=3-1\Rightarrow 2a=2\Rightarrow (2a)/2=2/2\Rightarrow a=1\) bit. Now something else that is important here: If the number is actually part of the solution, like \(a=1\) or \(a\le 1\) then you use a CLOSED circle on the line. If the number is above or below the solution, but not part of it, as in \(a < 1\) or \(a > 1\) then you use an OPEN circle. Filled in dot when it touches, hole when it does not. Keep that in mind as Nory gets you closer to the answer.

OpenStudy (anonymous):

@Hope_nicole

OpenStudy (anonymous):

you are basically turned it into sentence form...so when you read it, like how >, when you read it you say greater than, thats basically what you are doing

OpenStudy (anonymous):

so what do I do to get the answer?

OpenStudy (anonymous):

2x < -10 or x + 3>(or equal to) 11 A number line with a closed circle on -5, shading to the left, and an open circle on 8, shading to the right. A number line with an open circle on -5, a closed circle on 8, and shading in between. A number line with an open circle on -5, shading to the left, and a closed circle on 8, shading to the right. A number line with a closed circle on -5, an open circle on 8, and shading in between.

OpenStudy (anonymous):

so on a number line what does it look like?

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

wait no, B.

OpenStudy (anonymous):

like actually draw it silly!

OpenStudy (nory):

:)

OpenStudy (anonymous):

awwh:'( I'll try

OpenStudy (anonymous):

I can't do it, lol I keep messing up!

OpenStudy (anonymous):

just show me what you get!

OpenStudy (e.mccormick):